# Personalized Generation In Large Model Era: A Survey

### v1

@@ 3.2.3 Personalized Subjects
  • Encoder-based methods utilize a pre-trained image encoder to extract subject-specific features, which are then incorporated into text prompts or directly injected into the generator through dedicated cross-attention mechanisms or adapters. For instance, ELITE Wei et al. (2023), IP-Adapter Ye et al. (2023), Subject-Diffusion Ma et al. (2024a) and SSR-Encoder Zhang et al. (2024i) employ CLIP Radford et al. (2021) as the feature extractor, each adopting unique encoding and injection strategies to seamlessly incorporate subject features into the image generation process. MoMA Song et al. (2025) takes advantage of a pre-trained MLLM, LLaVA Liu et al. (2023b), to extract subject features and refine them based on the target prompt, producing contextualized image features that are injected into the generator via cross-attention layers. In addition, encoder-based methods have been extended to support multi-subject generation Patel et al. (2024b); Li et al. (2024d); Zhu et al. (2024); Wang et al. (2024i).
Furthermore, other research has explored diverse techniques for personalized subject-driven generation. These include prompt engineeringHe et al. (2024c), which formulates structured prompts to guide generation; instruction tuningZhou et al. (2024); Hu et al. (2024a), which fine-tunes models on personalized instructions for improved alignment; and reinforcement learning Chen et al. (2023d); Chae et al. (2023); Huang et al. (2024b); Wei et al. (2025a), which optimizes models through user feedback to refine subject representation.
@@ 3.2.4 Personal Face/Body
Generative models utilize personal face images to create high-fidelity portraits or avatars that preserve individual face IDs while adhering to users’ multimodal instructions, such as modifying expressions, actions, and backgrounds. Early GAN-based work mainly encodes face images into the latent space of StyleGAN Karras et al. (2019) for face manipulation Xia et al. (2021); Patashnik et al. (2021); Lyu et al. (2023); Baykal et al. (2023). To achieve more precise identity preservation and more flexible control, DM-based methods usually integrate a separate image encoder to convert face images into ID representations, which are then combined with user instructions to guide the face generation process. For instance, FastComposer Xiao et al. (2024b), Face0 Valevski et al. (2023), PhotoMaker Li et al. (2024k), Infinite-ID Wu et al. (2025a), MasterWeaver Wei et al. (2025b), and AddMe Yue et al. (2025) utilize a pre-trained CLIP image encoder or face recognition model to extract ID embeddings for identity preservation. In contrast, SeFi-IDE Li et al. (2024h) directly optimizes one ID representation as multiple per-stage tokens to enhance semantic control. Except for ID representations, some studies have explored additional features to enhance style control Yan et al. (2023), spatial control Wang et al. (2024g); He et al. (2024d, a); Jiang et al. (2025), especially specific human-object interaction  Guo et al. (2024b); Hu et al. (2025), and scene affordance Kulal et al. (2023); Parihar et al. (2024).
However, the rapid development of personalized face generation techniques has raised concerns about potential misuse and privacy risks. Recent research has investigated unlearning-related methods Wu et al. (2025b); Hu et al. (2024b), adversarial attack-based methods Van Le et al. (2023); Xiao et al. (2023); Wan et al. (2024); Onikubo and Matsui (2024); Liu et al. (2024e) and watermark-based methods Liu et al. (2024b) to protect user privacy.
@@
This task aims to synthesize a photorealistic image of a dress person by combining their body and face images with specified garments. Early GAN-based works Wang et al. (2018a); Dong et al. (2019a); Men et al. (2020); Choi et al. (2021); Lee et al. (2022) mainly follow a two-stage process. Initially, a dedicated warping module is employed to align garment images with the person’s body shape. Subsequently, the reshaped garment is seamlessly blended with the person’s image to generate the final try-on result. With the great success of DMs in various tasks, recent research has deployed their applications in virtual try-ons Zhang et al. (2024e); Ning et al. (2024); Kim et al. (2024a); Wan et al. (2025). For example, LaDI-VTON Morelli et al. (2023) and DCI-VTON Gou et al. (2023) explicitly warp clothes to match the person’s body and then utilize DMs for blending. In contrast, TryOnDiffusion Zhu et al. (2023a) proposes a Parallel-UNet architecture, which performs implicit warping and blending in a unified process. Similarly, several subsequent studies Choi et al. (2025); Xu et al. (2024d); Sun et al. (2024); Shen et al. (2024a) utilize parallel UNets for garment feature extraction and enhance blending via self-attention and cross-attention mechanisms.
@@ 3.2.5 Evaluation Metrics
To quantify the quality and coherence of generated images, conventional metrics such as Fréchet Inception Distance (FID) Heusel et al. (2017) and Kernel Inception Distance (KID) Bińkowski et al. (2018) are commonly employed. Beyond quantitative evaluation, most studies often include case studies and human evaluations to assess the personalization and instruction alignment of the generated images. In e-commerce scenarios, product image generation often incorporates online tests to evaluate model performance in real-world scenarios.
@@ 3.3.1 Personalized Subjects
Given the great success of various personalized models in subject-driven T2I generation, methods such as AnimateDiff Guo et al. (2024a), PIA Zhang et al. (2024h), and Still-Moving Chefer et al. (2024) adapt these models for T2V generation by incorporating motion and temporal dynamics through additional modules. MagDiff Zhao et al. (2025) enhances subject-driven video generation with three types of alignments. VideoBooth Jiang et al. (2024) proposes a coarse-to-fine manner to encode subject images into the generator. CustomCrafter Wu et al. (2024d) integrates a plug-and-play module with a dynamic weighted video sampling strategy to maintain motion generation and conceptual combination abilities during subject learning. Other studies introduce additional conditions, such as motion control Wu et al. (2024a); Wei et al. (2024b, c) and depth control He et al. (2023), enabling more flexible subject customization. Besides, some studies have explored multi-subject T2V generation Chen et al. (2023a, 2024b); Wang et al. (2024m). Beyond these, methods such as StyleCrafterLiu et al. (2024a) and StyleMaster Ye et al. (2024) integrate specified style images as subjects, allowing for stylized T2V generation that tailors the video aesthetic to the desired style.
@@ 3.3.2 Personal Face/Body
This task aims to synthesize lip-synchronized talking videos, typically driven by personal face images and corresponding audio clips. Recent research has explored diverse approaches, such as Neural Radiance Fields (NeRFs) Yao et al. (2022); Li et al. (2023b) and different backbone networks, including GANs Yi et al. (2020); Ki and Min (2023); Guan et al. (2023) and DMs Zhang et al. (2023b); Zhua et al. (2023); Liu et al. (2024f); Wei et al. (2024a); Tian et al. (2025); Tan et al. (2025); Wang et al. (2024d); Zheng et al. (2024b). Beyond audio-driven generation, some studies have also investigated video-driven Zhang et al. (2023a) and text-driven Choi et al. (2024) methods for talking head generation.
@@ 3.3.3 Evaluation Metrics
For video quality assessment, standard metrics include frame-level evaluations like FID and KID, as well as video-level metrics such as VFID Wang et al. (2018c), FID-VID Balaji et al. (2019), FVD Unterthiner et al. (2018), and KVD Unterthiner et al. (2018). Besides, some studies adopt CLIP-based cosine similarity between consecutive video frames to assess temporal consistency.
Beyond quantitative metrics, many works include qualitative case studies and human assessments to better evaluate personalization, instruction alignment, and overall video quality.
@@ 3.4.1 Personalized Subjects
Some methods further ensure alignment with textual prompts during the 3D generation process. MVDream Shi et al. (2023b) employs DMs to generate consistent multi-view images based on text prompts. DreamBooth3D Raj et al. (2023) combines DreamFusion and DreamBooth models within a three-stage optimization framework to enhance detail preservation and consistency through multi-view pseudo-data generation. Wonder3D Long et al. (2024) introduces a cross-domain diffusion model leveraging multi-view cross-attention to produce detailed normal and color maps. DreamFont3D Li et al. (2024e) utilizes NeRF as the 3D representation, optimizing font geometry and texture with multi-view mask constraints and progressive weight adjustments. Make-your-3D Liu et al. (2025) implements a joint optimization framework that combines identity-aware and subject-prior optimizations, aligning a 2D personalization model with a multi-view diffusion model for accurate 3D generation.
@@ 3.4.2 Personal Face/Body
In some cases, users may provide personal face and body images or monocular videos, aiming to generate identity-preserving 3D assets.
@@
For 3D face generation, Zhang et al. (2021a) introduces PoseGAN, a module designed to generate dynamic head poses. Gao et al. (2022) proposes a linear blend model based on multi-level voxel fields, representing expressions as neural radiance field bases. My3DGen Qi et al. (2023a) adapts the pre-trained EG3D model using Low-Rank Adaptation (LoRA) for parameter-efficient training on a limited set of personalized images. DiffusionTalker Chen et al. (2023c) employs contrastive learning to map speech features to personalized speaker identity. Wang et al. (2024h) integrates a face fusion module into a fine-tuned text-to-image diffusion model for identity-driven customization. Ko et al. (2024) utilizes VIVE3D to fine-tune the EG3D generator by inverting key frames from monocular videos. Song et al. (2024a) applies Cross-Modal Aggregation to blend style and facial motion features, ensuring alignment between generated facial expressions, speech, and styles. DiffSpeaker Ma et al. (2024d) proposes a diffusion model-based Transformer architecture to enhance speech-to-facial-animation mapping. In addition, FaceGPT Wang et al. (2024c) utilizes MLLMs for personalized 3D face understanding and generation from both user-provided face images and textual descriptions.
@@ 3.5.1 User behaviors
Methods like UMP Ma et al. (2022) and UP-Transformer Hu et al. (2022) infer user preferences by analyzing listening histories and ratings. UIGAN Wang et al. (2024k) adopts an interactive approach to collect user feedback, progressively refining the generated music. Ma et al. (2024b) construct a music community graph, where nodes represent users and songs, and edges capture relationships such as likes, subscriptions, and user interactions.
@@ 3.5.2 Personalized Subjects
Personalized text-to-audio generation explores methods for synthesizing tailored audio by aligning user preferences, textual descriptions, and contextual inputs. Mo et al. (2023) utilize Contrastive Language-Audio Pretraining (CLAP) to align features across audio and visual-text modalities in the spatiotemporal domain. Furthermore, Plitsis et al. (2024) adapt image-domain personalization techniques like Textual Inversion Gal et al. (2023) and DreamBooth Ruiz et al. (2023) to audio tasks, enhancing user-centric generation. Li et al. (2024j) introduce a Semantic-Aware Fusion (SAF) module to capture text-aware audio features, establishing nuanced perceptual relationships between text and audio inputs for more contextually aligned audio outputs.
@@ 3.5.3 Personal Face/Body
VioceMe van Rijn et al. (2022) employs SpeakerNet to derive speaker embeddings and incorporates Global Style Tokens (GST) for modeling speech styles. FR-PSS Wang et al. (2022) enhances the extraction of speech features from facial characteristics using a residual guidance strategy, which integrates prior speech information for improved efficiency. Lip2Speech Sheng et al. (2023) leverages a Variational Autoencoder (VAE) framework to disentangle speaker identity from linguistic content in videos, achieving fine-grained control over personalized speech synthesis via speaker embeddings.
@@ 3.6.2 User Documents
Several studies utilize user-created captions and comments to develop personalized captioning and comment systems Long et al. (2020); Zhang et al. (2020c); Geng et al. (2022). To effectively model user preferences, Shin et al. (2018) engages users by soliciting answers to targeted questions during generation. Lin et al. (2024b); Wu et al. (2024e) utilize users’ historical comments for personalized video comment generation. In addition, PV-LLM Lin et al. (2024b) fine-tunes MLLMs using user-written comments, while PVCG Wu et al. (2024e) learns a unique identifier for each user to enhance personalization.
@@ 3.6.3 Personalized subjects
  • Memory-based methods store user-specific subjects and activate or retrieve them as needed. For example, CSMN Chunseong Park et al. (2017); Park et al. (2018) stores image memory, user context memory, and word output memory, and the model generates personalized captions based on memory features. Hao et al. (2024) leverages RAG techniques to personalize LMMs, allowing LLMs to update their supported concepts without requiring additional training. PLVM Pham et al. (2024) proposes a pre-trained Aligner to align referential concepts with the queried images.
  • Optimization-based methods inject personalized information into the generation process via specific modules. For example, PerVL Cohen et al. (2022), Yo’LLaVA Nguyen et al. (2024b), and MC-LLaVA An et al. (2024) learn to encode a personalized subject into a set of latent tokens based on several provided subject images. MyVLM Alaluf et al. (2025) introduces learnable heads, each dedicated to recognizing a single user-specific subject. In addition, several studies have explored learning unique user embeddings Long et al. (2020); Zhang et al. (2020c); Xiong et al. (2020) or performing prefix tuning Wang et al. (2023f) for personalization.

### v2

@@ 3.2.3 Personalized Subjects
  • Encoder-based methods utilize a pre-trained image encoder to extract subject-specific features, which are then incorporated into text prompts or directly injected into the generator through dedicated cross-attention mechanisms or adapters. For instance, ELITE Wei et al. (2023), IP-Adapter Ye et al. (2023), Subject-Diffusion Ma et al. (2024a) and SSR-Encoder Zhang et al. (2024k) employ CLIP Radford et al. (2021) as the feature extractor, each adopting unique encoding and injection strategies to seamlessly incorporate subject features into the image generation process. MoMA Song et al. (2025) takes advantage of a pre-trained MLLM, LLaVA Liu et al. (2023b), to extract subject features and refine them based on the target prompt, producing contextualized image features that are injected into the generator via cross-attention layers. In addition, encoder-based methods have been extended to support multi-subject generation Patel et al. (2024b); Li et al. (2024d); Zhu et al. (2024); Wang et al. (2024i).
Furthermore, other research has explored diverse techniques for personalized subject-driven generation. These include prompt engineeringHe et al. (2024c), which formulates structured prompts to guide generation; instruction tuningZhou et al. (2024); Hu et al. (2024a), which fine-tunes models on personalized instructions for improved alignment; and reinforcement learning Chen et al. (2023e); Chae et al. (2023); Huang et al. (2024b); Wei et al. (2025a), which optimizes models through user feedback to refine subject representation.
@@ 3.2.4 Personal Face/Body
Generative models utilize personal face images to create high-fidelity portraits or avatars that preserve individual face IDs while adhering to users’ multimodal instructions, such as modifying expressions, actions, backgrounds, and styles. Early GAN-based work mainly encodes face images into the latent space of StyleGAN Karras et al. (2019) for face manipulation Xia et al. (2021); Patashnik et al. (2021); Lyu et al. (2023); Baykal et al. (2023). To achieve more precise identity preservation and more flexible control, DM-based methods usually integrate a separate image encoder to convert face images into ID representations, which are then combined with user instructions to guide the face generation process. For instance, FastComposer Xiao et al. (2024b), Face0 Valevski et al. (2023), PhotoMaker Li et al. (2024l), Infinite-ID Wu et al. (2025a), MasterWeaver Wei et al. (2025b), and AddMe Yue et al. (2025) utilize a pre-trained CLIP image encoder or face recognition model to extract ID embeddings for identity preservation. In contrast, SeFi-IDE Li et al. (2024h) directly optimizes one ID representation as multiple per-stage tokens to enhance semantic control. Except for ID representations, some studies have explored additional features or conditions to enhance style control Yan et al. (2023), spatial control Wang et al. (2024g); He et al. (2024d, a); Jiang et al. (2025), especially specific human-object interaction  Guo et al. (2024b); Hu et al. (2025), and scene affordance Kulal et al. (2023); Parihar et al. (2024).
However, the rapid development of personalized face generation techniques has raised concerns about potential misuse and privacy risks. Recent research has investigated unlearning-related methods Wu et al. (2025b); Hu et al. (2024b), adversarial attack-based methods Van Le et al. (2023); Xiao et al. (2023); Wan et al. (2024); Onikubo and Matsui (2024); Liu et al. (2024f) and watermark-based methods Liu et al. (2024b) to protect user privacy.
@@
This task aims to synthesize a photorealistic image of a dress person by combining their body and face images with specified garments. Early GAN-based works Wang et al. (2018a); Dong et al. (2019a); Men et al. (2020); Choi et al. (2021); Lee et al. (2022) mainly follow a two-stage process. Initially, a dedicated warping module is employed to align garment images with the person’s body shape. Subsequently, the reshaped garment is seamlessly blended with the person’s image to generate the final try-on result. With the great success of DMs in various tasks, recent research has deployed their applications in virtual try-ons Zhang et al. (2024e); Ning et al. (2024); Kim et al. (2024a); Wan et al. (2025). For example, LaDI-VTON Morelli et al. (2023) and DCI-VTON Gou et al. (2023) explicitly warp clothes to match the person’s body and then utilize DMs for blending. In contrast, TryOnDiffusion Zhu et al. (2023a) proposes a Parallel-UNet architecture, which performs implicit warping and blending in a unified process. Similarly, several subsequent studies Choi et al. (2025); Xu et al. (2024d); Sun et al. (2024a); Shen et al. (2024a) utilize parallel UNets for garment feature extraction and enhance blending via self-attention and cross-attention mechanisms.
@@ 3.2.5 Evaluation Metrics
To quantify the quality and coherence of generated images, conventional metrics such as Fréchet Inception Distance (FID) Heusel et al. (2017) and Kernel Inception Distance (KID) Bińkowski et al. (2018) are also commonly employed. Beyond quantitative evaluation, most studies present case studies and conduct human evaluations to assess the personalization and instruction alignment of the generated images. In e-commerce scenarios, product image generation often incorporates online tests to evaluate model performance in real-world scenarios.
@@ 3.3.1 Personalized Subjects
Given the great success of various personalized models in subject-driven T2I generation, methods such as AnimateDiff Guo et al. (2024a), PIA Zhang et al. (2024i), and Still-Moving Chefer et al. (2024) adapt these models for T2V generation by incorporating motion and temporal dynamics through additional modules. MagDiff Zhao et al. (2025) enhances subject-driven video generation with three types of alignments. VideoBooth Jiang et al. (2024b) proposes a coarse-to-fine manner to encode subject images into the generator. CustomCrafter Wu et al. (2024d) integrates a plug-and-play module with a dynamic weighted video sampling strategy to maintain motion generation and conceptual combination abilities during subject learning. Other studies introduce additional conditions, such as motion control Wu et al. (2024a); Wei et al. (2024b, c) and depth control He et al. (2023), enabling more flexible subject customization. Besides, some studies have explored multi-subject T2V generation Chen et al. (2023a, 2024b); Wang et al. (2024m). Beyond these, methods such as StyleCrafterLiu et al. (2024a) and StyleMaster Ye et al. (2024) integrate specified style images as subjects, allowing for stylized T2V generation that tailors the video aesthetic to the desired style.
@@ 3.3.2 Personal Face/Body
This task aims to synthesize lip-synchronized talking videos, typically driven by personal face images and corresponding audio clips. Recent research has explored diverse approaches, such as Neural Radiance Fields (NeRFs) Yao et al. (2022); Li et al. (2023b) and different backbone networks, including GANs Yi et al. (2020); Ki and Min (2023); Guan et al. (2023) and DMs Zhang et al. (2023b); Zhua et al. (2023); Shen et al. (2023); Liu et al. (2024g); Wei et al. (2024a); Tian et al. (2025); Tan et al. (2025); Wang et al. (2024d); Zheng et al. (2024b). Beyond audio-driven generation, some studies have also investigated video-driven Zhang et al. (2023a) and text-driven Choi et al. (2024) methods for talking head generation.
@@ 3.3.3 Evaluation Metrics
For overall video quality assessment, standard metrics include frame-level evaluations like FID and KID, as well as video-level metrics such as VFID Wang et al. (2018c), FID-VID Balaji et al. (2019), FVD Unterthiner et al. (2018), and KVD Unterthiner et al. (2018). Additionally, some studies leverage CLIP-based cosine similarity between consecutive video frames to assess temporal consistency.
Beyond quantitative metrics, many studies complement their evaluations with qualitative case studies and human assessments to better capture personalization, instruction alignment, and overall video quality.
@@ 3.4.1 Personalized Subjects
Some methods further ensure alignment with textual prompts during the 3D generation process. MVDream Shi et al. (2023b) employs a multi-view diffusion model to generate consistent multi-view images based on text prompts. DreamBooth3D Raj et al. (2023) combines DreamFusion and DreamBooth models within a three-stage optimization framework to enhance detail preservation and consistency through multi-view pseudo-data generation. Wonder3D Long et al. (2024) introduces a cross-domain diffusion model leveraging multi-view cross-attention to produce detailed normal and color maps. DreamFont3D Li et al. (2024e) utilizes NeRF as the 3D representation, optimizing font geometry and texture with multi-view mask constraints and progressive weight adjustments. Make-your-3D Liu et al. (2025a) implements a joint optimization framework that combines identity-aware and subject-prior optimizations, aligning a 2D personalization model with a multi-view diffusion model for accurate 3D generation.
@@ 3.4.2 Personal Face/Body
In some cases, users may provide personal face and body inputs in the form of images or monocular videos, aiming to generate identity-preserving 3D assets.
@@
For 3D face generation, Zhang et al. (2021a) introduces PoseGAN, a module designed to generate dynamic head poses. Gao et al. (2022) proposes a linear blend model based on multi-level voxel fields, representing expressions as neural radiance field bases. My3DGen Qi et al. (2023a) adapts the pre-trained EG3D model using Low-Rank Adaptation (LoRA) for parameter-efficient training on a limited set of personalized images. DiffusionTalker Chen et al. (2023d) employs contrastive learning to map speech features to personalized speaker identity. Wang et al. (2024h) integrates a face fusion module into a fine-tuned text-to-image diffusion model for identity-driven customization. Ko et al. (2024) utilizes VIVE3D to fine-tune the EG3D generator by inverting key frames from monocular videos. Song et al. (2024a) applies Cross-Modal Aggregation to blend style and facial motion features, ensuring alignment between generated facial expressions, speech, and styles. DiffSpeaker Ma et al. (2024d) proposes a diffusion model-based Transformer architecture to enhance speech-to-facial-animation mapping. Zero-1-to-A Zhou et al. (2025b) leverages pretrained video DMs to generate 4D avatars from a single face image, capturing both 3D spatial and temporal dynamics. In addition, FaceGPT Wang et al. (2024c) utilizes MLLMs for personalized 3D face understanding and generation from both user-provided face images and textual descriptions.
@@ 3.6.1 User behaviors
Methods like UMP Ma et al. (2022) and UP-Transformer Hu et al. (2022) infer user preferences by analyzing listening histories and ratings. UIGAN Wang et al. (2024k) adopts an interactive approach to collect user feedback, progressively refining the generated music. Ma et al. (2024b) constructs a music community graph, where nodes represent users and songs, and edges capture relationships such as likes, subscriptions, and user interactions. Furthermore, Plitsis et al. (2024) adapts image-domain personalization techniques like Textual Inversion Gal et al. (2023) and DreamBooth Ruiz et al. (2023) to audio tasks, enhancing user-centric music generation.
@@ 3.6.2 Personalized Subjects
Personalized text-to-audio generation explores techniques for synthesizing tailored audio by aligning user preferences, textual descriptions, and contextual inputs. YourTTS Casanova et al. (2022) focuses on zero-shot multi-speaker and multilingual training, allowing fast adaptation to unseen voices. Recent advancements such as Voicebox Le et al. (2023), StyleTTS 2 Li et al. (2023c), Make-An-Audio Huang et al. (2023), DiffAVA Mo et al. (2023), and TAS Li et al. (2024k), utilizes DMs for personalized text-to-audio generation. For instance, DiffAVA utilizes Contrastive Language-Audio Pretraining to align features across audio and visual-text modalities in the spatiotemporal domain. TAS introduces a Semantic-Aware Fusion module to capture text-aware audio features, establishing nuanced perceptual relationships between text and audio inputs for more contextually aligned audio outputs. Additionally, VALL-E X Zhang et al. (2023d) introduces a cross-lingual neural codec language model that supports personalized speech synthesis in a different language for monolingual speakers. Style-Talker Li et al. (2024i) builds a spoken dialogue system on a pretrained audio LLM, integrating user input audio, transcribed chat history, and speech styles to generate personalized audio responses.
@@ 3.6.3 Personal Face/Body
VioceMe van Rijn et al. (2022) employs SpeakerNet to derive speaker embeddings and incorporates Global Style Tokens for modeling speech styles. FR-PSS Wang et al. (2022a) enhances the extraction of speech features from facial characteristics using a residual guidance strategy, which integrates prior speech information for improved efficiency. Lip2Speech Sheng et al. (2023) leverages a Variational Autoencoder framework to disentangle speaker identity from linguistic content in videos, achieving fine-grained control over personalized speech synthesis via speaker embeddings.
@@ 3.7.2 User Documents
Several studies utilize user-created captions and comments to develop personalized captioning and comment systems Shuster et al. (2019); Long et al. (2020); Zhang et al. (2020c); Geng et al. (2022). To effectively model user preferences, Shin et al. (2018) engages users by soliciting answers to targeted questions during generation. Lin et al. (2024b) and Wu et al. (2024e) utilize users’ historical comments for personalized video comment generation. SimTube Hung et al. (2024) incorporates user responses and defined personas to support user-steerable video comment generation.
In addition, PV-LLM Lin et al. (2024b) fine-tunes MLLMs using user-written comments, while PVCG Wu et al. (2024e) learns a unique identifier for each user to enhance personalization.
@@ 3.7.3 Personalized subjects
  • Memory-based methods store user-specific subjects and activate or retrieve them as needed. For example, CSMN Chunseong Park et al. (2017); Park et al. (2018) stores image memory, user context memory, and word output memory, and the model generates personalized captions based on memory features. Hao et al. (2024) and Seifi et al. (2025) leverage RAG techniques to personalize pretrained MLLMs, allowing LLMs to update their supported concepts without requiring additional training. PLVM Pham et al. (2024) proposes a pre-trained Aligner to align referential concepts with the queried images. In addition, R2P Das et al. (2025) introduces a retrieval-reasoning paradigm to identify personal concepts in a training-free setting.
  • Optimization-based methods inject personalized information into the generation process via specific modules. For example, PerVL Cohen et al. (2022), Yo’LLaVA Nguyen et al. (2024b), and MC-LLaVA An et al. (2024) learn to encode a personalized subject into a set of latent tokens based on several provided subject images. MyVLM Alaluf et al. (2025) introduces learnable heads, each dedicated to recognizing a single user-specific subject. CaT An et al. (2025) proposes a comprehensive data synthesis pipeline to enhance the personalization capabilities of MLLMs. PVChat Shi et al. (2025) empowers LLMs with personalization capabilities for subject-aware question answering with only one-shot learning. In addition, several studies have explored learning unique user embeddings Long et al. (2020); Zhang et al. (2020c); Xiong et al. (2020) or performing prefix tuning Wang et al. (2023f) for personalization.